Can someone help me on this problem?
I need to solve for x and check results:
6x - 1/5 - 2x/3 = 3
Thanks,
Sher

Answer by Juhi(44)

6x - 1/5 - 2x/3 = 3
6x - 2x/3 = 3 + 1/5

16x/3 = 16x/5
x = 16x/5 * 3/16x
x = 3/5